OpenStudy (anonymous):

log(down)4 (x^2-3x)=1

OpenStudy (anonymous):

\[\log_{4} (x ^{2}-3x)=1\]

jimthompson5910 (jim_thompson5910):

\[\log_{4}(x^{2}-3x)=1\] \[x^{2}-3x=4^1\] \[x^{2}-3x=4\] \[x^{2}-3x-4=0\] \[(x-4)(x+1)=0\] \[x-4=0 \ \text{or} \ x+1=0\] \[x=4 \ \text{or} \ x=-1\]

OpenStudy (anonymous):

THank you!!

jimthompson5910 (jim_thompson5910):

yw